OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x- 3y- 5 = 0 2x+ 3y- 5 = 0 -2x+ 3y- 5 = 0

OpenStudy (mertsj):

First find the slope.

OpenStudy (anonymous):

As an investment in your math career, download GeoGebra. Its free! Your answer is B, 2x+3y-5=0

OpenStudy (anonymous):

Find the slope and substitute into eq:\[y = mx + b\] where m is the slope you will first calculate. Using any one pair of points that you r given just pulg it into the eqquation, you eill get the "b" value. Then you will have the equation.

OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 7) and (4, 9)? slope = -1/3 slope = 1/3 slope = 3